OpenStudy (anonymous):

Write the following equation set equal to zero and in factored form. (x - 2)(x - 3) = 2 (x - 1)(x - 4) = 0 (x - 2)(x - 3) = 0 (x - 4)(x - 5) = 0

OpenStudy (anonymous):

\[x ^{2}-5x+6-2 = x ^{2}-5x+4 = (x-1)(x-4)=0\] (x-1)(x-4)=0